OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −1/5(x − 2) y − 2 = −1/5(x + 3)

OpenStudy (anonymous):

A line that has those points must have what slope?

OpenStudy (anonymous):

slope = 1-2/(2+3) = -1/5 hence answer is d